# Towards verifications of Krylov complexity

### v1

@@ 6.2 Hermite and Laguerre
These two are the best known examples of exactly solvable one-dimensional quantum mechanical systems, the harmonic oscillator for the Hermite (H) and the harmonic oscillator with a centrifugal potential for the Laguerre (L). Like K and dH cases in §5.1, all the moments form geometrical sequences, $\mu_{2m}=\lambda^{2(m-1)}\mu_{2}$, $\lambda=2$ for H and $\lambda=4$ for L. The Krylov orthogonalisation stops at $\mathcal{O}_{2}$. The data for each system are:
@@
As mentioned above, the Heisenberg operator solution for $x^{2}+1/x^{2}$ (L) and $1/\sin^{2}x$ (G) potentials are reported by Nieto and Simmons [19].
@@ 7 Summary and comments
For a group of exactly solvable compact discrete quantum systems, the moments of the operators in a Krylov subspace spanned by a Liouville operator $\mathcal{L}:=[\mathcal{H},\cdot]$ and the sinusoidal coordinate $\eta$ are evaluated explicitly. They provide the essential tool for measuring the growth of operators evolving under Hamiltonian dynamics [1]. Understanding of the complexity of exactly solvable quantum systems would reveal the nature of quantum chaos by contrast. The moments of exactly solvable non-compact discrete dynamics and ordinary one-dimensional quantum systems are also evaluated explicitly by adopting the inner product involving the Boltzmann factor. These exactly solvable discrete quantum mechanical systems can be regarded as a very special type of matrix models.
@@
Four explicit examples of multivariate discrete orthogonal polynomials, the multivariate Krawtchouk, Meixner and two types of Rahman polynomials, are constructed recently by myself [22]. They are eigenvectors of respective Hamiltonians having nearest neighbour interactions. Investigation of their complexity through Krylov orthonormalisation would expose the contrast between integrable and chaotic multi-particle dynamics.
@@ Acknowledgements
R. S. thanks Jen-Chi Lee for the hospitality and vigorous discussions during his visit to National Yang Ming Chiao Tung University. He also thanks Yi Yang for opening his eyes to the topic of Krylov complexity.

### v2

@@ 6.2 Hermite and Laguerre
These two are the best known examples of exactly solvable one-dimensional quantum mechanical systems, the harmonic oscillator for the Hermite (H) and the harmonic oscillator with a centrifugal potential for the Laguerre (L). Like K and dH cases in §5.1, all the moments form geometrical sequences, $\mu_{2m}=\lambda^{2(m-1)}\mu_{2}$, $\lambda=2$ for H and $\lambda=4$ for L. The Krylov orthogonalisation stops at $\mathcal{O}_{2}$, a clear sign of ‘non-complexity’ of these solvable systems. The data for each system are:
@@
As mentioned above, the Heisenberg operator solution for $x^{2}+1/x^{2}$ (L) and $1/\sin^{2}x$ (G) potentials are reported by Nieto and Simmons [21][24].
@@ 7 Summary and comments
For a group of exactly solvable compact discrete quantum systems, the moments of the operators in a Krylov subspace spanned by a Liouville operator $\mathcal{L}:=[\mathcal{H},\cdot]$ and the sinusoidal coordinate $\eta$ are evaluated explicitly. They provide the essential tool for measuring the growth of operators evolving under Hamiltonian dynamics [1]. Understanding of the complexity of exactly solvable quantum systems would reveal the nature of quantum chaos by contrast. The moments of exactly solvable non-compact discrete dynamics and ordinary one-dimensional quantum systems are also evaluated explicitly by adopting Wightman inner product involving the Boltzmann factor. These exactly solvable discrete quantum mechanical systems can be regarded as a very special type of matrix models.
@@
Four explicit examples of multivariate discrete orthogonal polynomials, the multivariate Krawtchouk, Meixner and two types of Rahman polynomials, are constructed recently by myself [27][29]. They are eigenvectors of respective Hamiltonians having nearest neighbour and other types of interactions. Investigation of their complexity through Krylov orthonormalisation would expose the contrast between integrable and chaotic multi-particle dynamics.
@@ Acknowledgements
R. S. thanks Jen-Chi Lee for the hospitality and vigorous discussions during his visit to National Yang Ming Chiao Tung University. He also thanks Yi Yang for opening his eyes to the topic of Krylov complexity. Many useful suggestions by referees are greatly acknowledged.
Declarations
  • Funding: No funds, grants, or other support was received.
  • Data availability statement: Data sharing not applicable to this article as no datasets were generated or analysed during the current study.
  • Competing Interests: The author has no competing interests to declare that are relevant to the content of this article.